# Social Science Is Necessary for Operationalizing Socially Responsible Foundation Models

### v1

@@ 4 Facilitating Socially Responsible Foundation Model Research
  • The evaluation criteria for social scientists’ careers may place less value on (1) AI-area or interdisciplinary AI + social science venues, or (2) conference publications compared to journal publications, which could discourage them from publishing in top-tier AI venues; and vice-versa for AI researchers. Like the FAccT conference,444 See https://facctconference.org/2024/cfp. more AI conferences could offer the optional choice of non-archival paper submissions (in addition to the standard archival submission), allowing researchers from other fields to later submit their conference papers to discipline-specific journals.
@@
  • To incentivize research on interdisciplinary AI + social science topics, institutions could better consider interdisciplinary work in career advancement [77] and funding proposals assessment [78], and offer specialized funding opportunities and sabbaticals, allowing researchers to explore new ideas and collaborations in a wider context [79].
@@
  • Promoting interdisciplinary education can better prepare the next generation of researchers – e.g., awarding degree credit for courses taken in other fields encourages students to learn the essentials of other disciplines [80], providing them with the necessary foundations to integrate methods from, and facilitate collaborate with, fields beyond their primary research area.
@@
  • Researchers who wish to embrace a more interdisciplinary agenda could broaden their expertise with workshops, tutorials, or short courses provided by researchers from other fields. For example, we recommend AI conferences to open tutorial calls to non-AI researchers; and likewise for social science venues.
@@ 5 Conclusion
In this work, we have advocated for interdisciplinary research between AI and social science in the context of foundation models like LLMs, focusing on the importance of social science in understanding the affordances and social impacts of such transformative technologies. We outlined the importance of interdisciplinary expertise and collaboration throughout the foundation model R&D pipeline, highlighted the associated responsibilities and benefits for model providers and deployers, and provided actionable suggestions to promote collaboration between AI and social science.
Takeaways
We recommend that AI experts and labs researching, developing, or deploying foundation models reflect on incorporating interdisciplinary collaboration within their team and their research topic more broadly, particularly in promoting socially responsible affordances and studying potential social impacts of their work. Neither AI nor social science holds all the answers regarding how to develop safe, beneficial, and socially responsible foundation models; and it is critical that both disciplines work more closely together toward this goal, rather than “siloing” research for such a potentially transformative technology.
Limitations and Future Work
The primary limitation of this work is that interdisciplinary research can be expensive and time-consuming [77], and bringing in diverse perspectives always carries the potential to dilute research focus with competing visions and priorities [81]. In Section 4, we outline a few tentative suggestions to reduce this cost; but they are by no means comprehensive. We recommend that future work consider performing more comprehensive cost-benefit analyses along multiple dimensions (cf. [82]) to assess the resources needed to achieve the benefits outlined above, making it possible to more effectively manage these research tradeoffs.
@@ Acknowledgements
We thank Alan Craig, Dave Buckley, and Linda Derhak for their help in facilitating this project and sharing valuable feedback on an earlier version of this draft. This work is supported in part by the National Science Foundation and the Institute of Education Sciences, U.S. Department of Education, through Award #2229612 (National AI Institute for Inclusive Intelligent Technologies for Education). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of National Science Foundation or the U.S. Department of Education. This material is based upon work supported in part by the National Science Foundation under Grant No. 2217706. The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Elisa Nguyen. This work was supported in part by the Tübingen AI Center and the Parameter Lab company.

### v2

@@ 4 Facilitating Socially Responsible Foundation Models
  • As in the FAccT conference,444 See https://facctconference.org/2024/cfp. more AI conferences could offer the optional choice of non-archival paper submissions (in addition to the standard archival submission), allowing researchers from other fields to later submit their conference papers to discipline-specific journals.
@@
  • Research institutions could better consider interdisciplinary work in career advancement Pellmar et al. (2000) and funding proposals assessment Bromham et al. (2016), and offer specialized funding opportunities and sabbaticals, allowing researchers to explore new ideas and collaborations in a wider context Ioppolo & Wooding (2023).
  • Existing practices for promoting socially-responsible research can be further promoted by publication venues (e.g., by making model and dataset cards Mitchell et al., 2019 a mandatory part of certain submission types) to expand their adoption as a standard in the research community.
@@
  • Promoting interdisciplinary education helps provide the next generation of researchers with the foundations to integrate methods from, and facilitate collaborations with, fields beyond their primary research area. For instance, awarding degree credit for courses from other disciplines encourages students to learn the essentials of these fields Amelink et al. (2024), and embedding ethics education in technical courses can improve students’ abilities to engage in relevant ethical discussions Horton et al. (2022).
@@
  • Researchers considering a more interdisciplinary agenda could broaden their expertise with workshops, tutorials, or short courses provided by researchers from other fields. For example, we suggest that AI and social science conferences open tutorial calls to researchers outside their respective disciplines.
@@ 5 Conclusion
In this work, we have advocated for interdisciplinary research between AI and social science in the context of foundation models like LLMs, focusing on the importance of social science in understanding the affordances and social impacts of such transformative technologies. We outlined the importance of interdisciplinary expertise and collaboration throughout the foundation model R&D pipeline, highlighted the associated responsibilities and benefits for model providers and deployers, and provided actionable suggestions to promote collaboration between AI and social science. Finally, we discuss a few important considerations for future work in Appendix A.
@@ Acknowledgements
We thank Alan Craig, Dave Buckley, and Linda Derhak for their help in facilitating this project and sharing valuable feedback. This work is supported in part by the National Science Foundation and the Institute of Education Sciences, U.S. Department of Education, through Award #2229612 (National AI Institute for Inclusive Intelligent Technologies for Education). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of National Science Foundation or the U.S. Department of Education. This material is based upon work supported in part by the National Science Foundation under Grant No. 2217706. The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Elisa Nguyen. This work was supported in part by the Tübingen AI Center and the Parameter Lab company.
Appendix A Future Work
An important consideration regarding the interventions suggested in Section 4 is that interdisciplinary research can be expensive and time-consuming Pellmar et al. (2000), and bringing in diverse perspectives always carries the potential to dilute research focus with competing visions and priorities Committee on Science and Public Policy and Committee on Facilitating Interdisciplinary Research (2005). We suggest that future work could consider performing more comprehensive cost-benefit analyses along multiple dimensions (cf. Slaper et al., 2011) to assess the resources needed to achieve the benefits outlined above, making it possible to more effectively manage these research tradeoffs. More broadly, we recommend that AI experts and labs researching, developing, or deploying foundation models reflect on incorporating interdisciplinary collaboration within their team and their research topic more broadly, particularly in promoting socially responsible affordances and studying potential social impacts of their work. Neither AI nor social science holds all the answers regarding how to develop safe, beneficial, and socially responsible foundation models; and it is critical that both disciplines work more closely together toward this goal, rather than “siloing” research for such a potentially transformative technology.